OpenStudy (anonymous):

Solve the equation. 2|2x – 12| + 18 = 54

OpenStudy (xapproachesinfinity):

subtract 18 from both side, then divide both side by 2

OpenStudy (bradely):

2|2x – 12| + 18-18 = 54-18 2|2x – 12| = 36 |2x – 12| = 18 2x-12=18 or 2x – 12=-18 2x=30 2x=-6 x=15 x=-3 Source: http://www.mathskey.com/question2answer/